OpenStudy (anonymous):

Did I do anything wrong? 1/2b-2=1 1/2b*2/1-2=1 b=2=1 b-2+2=1+2 b=3 Will give medal!!!

OpenStudy (anonymous):

It would be easier to do this \[\frac{ 1 }{ 2 }b - 2 = 1\] \[\frac{ 1 }{ 2 }b = -1\] Then multiply both sides by (1/2). What's the answer?